Moa invested $80000,part at 4% and the rest at 2%.Find the amount invested at each rate if the annual income from the two investmets is $2500.
Answer by mananth(16946)

x+y80000 ------------------------1
4.00%x+ 2.00%y=$2,500.00
Multiply by 100
4x+2y= $250,000.00 --------2

Multiply (1) by -4
we get

-4x-4y= -320000.00

Add this to (2)

-2y= -$70,000.00

divide by -2

y=$35,000.00 investment at 2.00%
Balance $45,000.00 investment at 4.00%
CHECK Interest
$45,000.00 $1,800.00
$35,000.00 $700.00
Total $2,500.00